1. Solve the system of equations: {y=3x-14}
{ y=5x-20 }



Answer by checkley75(3666)

Y=3X-14
Y=5X-20 SEEING AS EACH IS EQUAL TO Y THEN THEY ARE EQUAL TO EACH OTHER tHEN:
3X-14=5X-20
3X-5X=-20+14
-2X=-6
X=-6/-2
X=3 ANSWER.
Y=3*3-14
Y=9-14
Y=-5 ANSWER.
PROOF
-5=5*3-20
-5=15-20
-5=-5